Hello, I was helping my aunt with some of the problems she had with inequalities before i stumbled upon this. Its been years since I last took algebra, so please help me out with this.
1. In 2002, a one year old Acura was worth $18,249, with a depreciation of $1903 per year. A year old Subaru sedan was worth about $14,564 with a depreciation of $1225 per year.
(a) Let V be the value in $ of the Acura at t years since 2002, find an equation of a model to describe the data.
(b) Let V be the value in $ of the Subaru car at t years since 2002, find an equation of a model to describe the data.
(c) Use substitution or elimination to predict when the two cars will have the same value. What is that value.
Thanks.

Answer by scott8148(6628)

(a) V = 18249 - 1903t

(b) V = 14564 - 1225t

(c) substituting ___ 18249 - 1903t = 14564 - 1225t

adding 1903t-14564 ___ 3685 = 678t

dividing by 678 ___ t = 3685 / 678 = 5.435 (approx) ___ about 5 1/2 yrs from 2002

substituting ___ V = 18249 - 1903 (3685 / 678) = 7906